# Dynamic Context Compression for Efficient RAG

### v2

@@ Abstract
Retrieval-augmented generation (RAG) enhances large language models (LLMs) with external knowledge but incurs significant inference costs due to lengthy retrieved contexts. While context compression mitigates this issue, existing methods apply fixed compression rates—over-compressing simple queries or under-compressing complex ones. We propose Adaptive Context Compression for RAG (ACC-RAG), a framework that dynamically adjusts compression rates based on input complexity, optimizing inference efficiency without sacrificing accuracy. ACC-RAG combines a hierarchical compressor (for multi-granular embeddings) with a context selector to retain minimal sufficient information, akin to human skimming. Evaluated on Wikipedia and five QA datasets, ACC-RAG outperforms fixed-rate methods and matches/unlocks >4× faster inference versus standard RAG while maintaining or improving accuracy.
Dynamic Context Compression for Efficient RAG
Shuyu Guo Shandong University Qingdao, China guoshuyu225@gmail.com          Zhaochun Ren Leiden University Leiden, The Netherlands z.ren@liacs.leidenuniv.nl
@@ 1 Introduction
Large Language Models (LLMs) are pre-trained on massive datasets, encoding vast knowledge within billions of parameters. While they excel at many tasks, their parametric knowledge often falls short for knowledge-intensive applications. Retrieval-Augmented Generation (RAG) addresses this limitation by extending the model’s knowledge boundaries through external context retrieval Gao et al. (2023). However, integrating lengthy retrieved contexts into prompts increases inference costs and may exceed LLMs’ context window limits Chevalier et al. (2023)2.
@@
Figure 1: A case of the retrieval augmented model with different context compression methods.
@@
In summary, our contributions include: (1) A novel adaptive context compression framework for RAG, improving inference efficiency while maintaining retrieval augmentation benefits. (2) Training strategies for a hierarchical compressor and adaptive selector, enabling multi-granularity compressed embedding generation and automatic context selection. (3) A unified benchmark for context compression methods, resolving evaluation biases caused by inconsistent training data and task setups in prior works. (4) Experimental results showing our method as the best under context compression, with comparable results to regular RAG but with improved efficiency. (5) Further analysis highlighting the significant potential of adaptive context compression.
@@ 4.1 Datasets
To ensure a fair comparison of different compression methods, we train and evaluate all methods on unified datasets. For the retrieval corpus, we use the Wikipedia corpus from Dec. 20, 2018, standardized by Karpukhin et al. (2020) using the preprocessing from Chen et al. (2017). The corpus is split into multiple, disjoint text blocks of 128 tokens as documents, resulting in 21,015,324 documents in the end. For finetuning, we conduct experiments on five commonly used open-domain QA benchmarks, namely Natural Questions (NQ) Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), WebQuestions (WQ) Berant et al. (2013), CuratedTREC (TREC) Baudis and Sedivý (2015), and SQuAD v1.1 Rajpurkar et al. (2016). All question-answer pairs are ensured to have supporting documents available in the retrieval corpus. For pretraining, we select all supporting documents from the training set of QA and randomly sample additional documents from the retrieval corpus, resulting in a total of 1 million documents for training.
@@ 4.2 Baselines
xRAG Cheng et al. (2024): Maps retrieval embeddings into compression embeddings for extreme compression. Performs only auto-encoding during pretraining and combines instruction-following with self-distillation during finetuning.
@@ 6 Conclusion
In this paper, we propose Adaptive Context Compression for RAG, a framework that dynamically adjusts compressed embeddings required during inference. The framework effectively improves inference efficiency while maintaining quality benefit from retrieval. Experimental results under standardized benchmark show that our method achieves the best results under context compression and comparable or superior accuracy with over 4× inference efficiency compared to RAG.
@@
We acknowledge the importance of the ACM code of Ethics and totally agree with it. We ensure that this work is compatible with the provided code. All data used in this study is obtained from existing benchmarks, while all models used are publicly available, thus ensuring a high level of transparency and reproducibility in our experimental procedure.
@@ Appendix A Datasets
We use the dataset version processed by DPR [^1] https://github.com/facebookresearch/DPR including retrieval corpus and five QA datasets, which is licensed by CC-BY-NC 4.0. All datasets are English-language resources within the question answering (QA) domain. The datasets we used inherit from them while maintaining consistency in QA intent. We ensure that the dataset is free from harmful content or information leakage through a sampling approach. The detailed statistics of the dataset are presented in Table 5. Notably, based on whether the supporting documents are present in the pretraining dataset, the development set of the finetuning data is further divided into seen and unseen subsets.
@@ Appendix C Implementation Details
In Table 6 and Table 7, we list the hyperparameters for compressor pretraining and finetuning. These parameters are primarily inherited from prior work (e.g., learning rate, LoRA configuration), as the same backbone LLM and similar training tasks are used. Additionally, all methods are trained under the same configuration for a fair comparison, ensuring no intentional optimization bias towards our approach. For efficient training, at each training step, we randomly sample only one granularity from granularity sequence and one task to optimize during pretraining. In our configuration, we pretrain the compressor with roughly 20 hours and finetune with roughly 50 hours. For selector training, we list the hyperparameters in Table 8. We primarily search the hyperparameters of learning rates and find that a range between $2e^{-4}$ and $1e^{-4}$ yields the best results. Additionally, we adjusted the number of attention heads and layers in the transformer encoder. We observed that increasing both parameters beyond 4 provided no improvement, while reducing them below 4 resulted in a slight decline in performance. The entire training process took approximately 1 hour, and we ultimately selected the model with the highest sequence accuracy on the test set. All experiments are implemented using PyTorch [^6] https://pytorch.org/ and Hugging Face Transformers [^7] https://github.com/huggingface/transformers.
@@
Table 7: Hyperparameters for compressor finetuning.
@@ Appendix E Analysis on Granularity Sequences of Compressor
The results of different granularity sequences for compressor under inference granularity sequence $[1,2,8,32]$ and $[1,32]$ are shown in Figure 6 and Figure 7. The conclusion are consistent with Section 5.2.3. We observe that the final accumulated M scores exceeds the vanilla RAG even with only two inference granularities.
@@ Appendix I OOD Analysis
Generated with 0 context tokens, wrong answer.

### v3

# Enhancing RAG Efficiency with Adaptive Context Compression

@@ Abstract
Retrieval-augmented generation (RAG) enhances large language models (LLMs) with external knowledge but incurs significant inference costs due to lengthy retrieved contexts. While context compression mitigates this issue, existing methods apply fixed compression rates—over-compressing simple queries or under-compressing complex ones. We propose Adaptive Context Compression for RAG (ACC-RAG), a framework that dynamically adjusts compression rates based on input complexity, optimizing inference efficiency without loss of accuracy. ACC-RAG combines a hierarchical compressor (for multi-granular embeddings) with a context selector to retain minimal sufficient information, akin to human skimming. Evaluated on Wikipedia and five QA datasets, ACC-RAG outperforms fixed-rate methods and unlocks >4× faster inference versus standard RAG while maintaining or improving accuracy.
Enhancing RAG Efficiency with Adaptive Context Compression
Shuyu Guothanks: Work performed during a visit to Leiden University. Shandong University Qingdao, China guoshuyu225@gmail.com          Shuo Zhang Bloomberg London, United Kingdom szhang611@bloomberg.net          Zhaochun Renthanks: Corresponding Author. Leiden University Leiden, The Netherlands z.ren@liacs.leidenuniv.nl
@@ 1 Introduction
Large Language Models (LLMs) are pre-trained on massive datasets, encoding vast knowledge within billions of parameters. While they excel at many tasks, their parametric knowledge often falls short for knowledge-intensive applications. Retrieval-Augmented Generation (RAG) addresses this limitation by extending the model’s knowledge boundaries through external context retrieval Gao et al. (2023). However, integrating lengthy retrieved contexts into prompts increases inference costs and may exceed LLMs’ context window limits Chevalier et al. (2023).
@@
Figure 1: An example of a retrieval-augmented model with different context compression methods.
@@
In summary, our contributions include: (1) A novel adaptive context compression framework for RAG, improving inference efficiency while maintaining retrieval augmentation benefits. (2) Training strategies for a hierarchical compressor and adaptive selector, enabling multi-granularity compressed embedding generation and automatic context selection. (3) A unified benchmark for context compression methods, resolving evaluation biases from inconsistent training data and task setups in prior works. (4) Experimental results showing our method as the best under context compression, with comparable results to regular RAG but with improved efficiency. (5) Further analysis highlighting the potential of adaptive context compression.
@@ 4.1 Datasets
To ensure a fair comparison of different compression methods, we train and evaluate all methods on unified datasets. For the retrieval corpus, we use the Wikipedia corpus from Dec. 20, 2018, standardized by Karpukhin et al. (2020) using the preprocessing from Chen et al. (2017). The corpus is split into multiple, disjoint text blocks of 128 tokens as documents, resulting in 21,015,324 documents in the end. For fine-tuning, we conduct experiments on five commonly used open-domain QA benchmarks, namely Natural Questions (NQ) Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), WebQuestions (WQ) Berant et al. (2013), CuratedTREC (TREC) Baudis and Sedivý (2015), and SQuAD v1.1 Rajpurkar et al. (2016). All question-answer pairs are ensured to have supporting documents available in the retrieval corpus. For pretraining, we select all supporting documents from the training set of QA and randomly sample additional documents from the retrieval corpus, resulting in a total of 1 million documents for training.
@@ 4.2 Baselines
xRAG Cheng et al. (2024): Maps retrieval embeddings into compression embeddings for extreme compression. Performs only auto-encoding during pretraining and combines instruction-following with self-distillation during fine-tuning.
@@ 6 Conclusion
In this paper, we propose Adaptive Context Compression for RAG, a framework that dynamically adjusts compressed embeddings required during inference. The framework effectively improves inference efficiency while maintaining the quality benefits of retrieval. Experimental results under standardized benchmark show that our method achieves the best results under context compression and comparable or superior accuracy with over 4× inference efficiency compared to RAG.
@@
We acknowledge the importance of the ACM Code of Ethics and affirm our adherence to it. We ensure that this work is compatible with the provided code. All data used in this study is obtained from existing benchmarks, while all models used are publicly available, thus ensuring a high level of transparency and reproducibility in our experimental procedure.
@@ Appendix A Datasets
We use the dataset version processed by DPR [^1] https://github.com/facebookresearch/DPR including retrieval corpus and five QA datasets, which is licensed by CC-BY-NC 4.0. All datasets are English-language resources within the question answering (QA) domain. The datasets we used inherit from them while maintaining consistency in QA intent. We ensure that the dataset is free from harmful content or information leakage through a sampling approach. The detailed statistics of the dataset are presented in Table 5. Notably, based on whether the supporting documents are present in the pretraining dataset, the development set of the fine-tuning data is further divided into seen and unseen subsets.
@@ Appendix C Implementation Details
In Table 6 and Table 7, we list the hyperparameters for compressor pretraining and fine-tuning. These parameters are primarily inherited from prior work (e.g., learning rate, LoRA configuration), as the same backbone LLM and similar training tasks are used. Additionally, all methods are trained under the same configuration for a fair comparison, ensuring no intentional optimization bias towards our approach. For efficient training, at each training step, we randomly sample only one granularity from granularity sequence and one task to optimize during pretraining. In our configuration, we pretrain the compressor with roughly 20 hours and finetune with roughly 50 hours. For selector training, we list the hyperparameters in Table 8. We primarily search the hyperparameters of learning rates and find that a range between $2e^{-4}$ and $1e^{-4}$ yields the best results. Additionally, we adjusted the number of attention heads and layers in the transformer encoder. We observed that increasing both parameters beyond 4 provided no improvement, while reducing them below 4 resulted in a slight decline in performance. The entire training process took approximately 1 hour, and we ultimately selected the model with the highest sequence accuracy on the test set. All experiments are implemented using PyTorch [^6] https://pytorch.org/ and Hugging Face Transformers [^7] https://github.com/huggingface/transformers.
@@
Table 7: Hyperparameters for compressor fine-tuning.
@@ Appendix E Analysis on Granularity Sequences of Compressor
The results of different granularity sequences for compressor under inference granularity sequence $[1,2,8,32]$ and $[1,32]$ are shown in Figure 6 and Figure 7. The conclusions are consistent with Section 5.2.3. We observe that the final accumulated M scores exceeds the vanilla RAG even with only two inference granularities.
@@ Appendix I OOD Analysis
Generated with 0 context tokens (incorrect answer).